From PREX Wiki
Jump to: navigation, search

Back to Main Page >> Analyzer_Meeting

previous meeting << >> following meeting

Logistic information

Meeting room at JLab: CC L210A (entrance is in the lobby across from the 2nd floor auditorium entrance)

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  419 045 306
 Bluejeans link: https://bluejeans.com/419045306


  1. runlist
    • What do we need to do before we're ready to respin all of the concurrent running
      • Look at the event cuts---Tao
        • TY: will check the number of events cut around a low/high beam cut and after I'll replay the 1400s run#s.
      • SAM pedestals---Tao and Devi (Tao has found different beam-dependent pedestals, should we use those or just the beam-off peds?)
        • TY: for now we'll go with the pedestals we have now.
      • Do we need to mess with beamline pedestals? Not for this pass.
      • We'd like to have a run-summary tree: CC: working on it and should be done soon. CG: we'll test your code on the runs Tao replays and go from there. Expect presentation on Friday.
      • We want to have the post-pan regression feeding in into the run-summary tree (or friend-able) CC: functionality will be there from the start for the summary tree
      • SR: will have some results for the text output soon (late this week or next week)
  2. "Operations" branch
    • PK: we should talk to Caryn and figure out if she changed the configuration in the CH crate so that we update it on the japan channel mapping.
  3. Pull requests
    • PK: merged dev->master->operations
    • PK: Tao is cleaning up the beam line channel mapping for the APEX running. Will coordinate with Tao on how we should proceed.
  4. online plots
    • PK: should work now. Closing this issue.
  5. pair analysis
    • PK: Tao is working on the pair tree output.
  6. regression analysis
    • RR: we are transitioning from hard coded variables to arrays. Should have all the data handler functionality done by mid April
  7. feedback
    • AP: trying to do the PITA feedback, position feedback (should be done by Friday). IAx3 and helicity magnets still need to be implemented.
  8. Green Monster: there needs to be ways to set all the VQWKs/timing boards/etc consistently.
    • AZ: have a bug in the pyGM and working on it.
    • AZ: Kent wanted to have setting in the new GM to set values on the integration of the VQKs
  9. VQWK adcs
    • CC: one channel still needs to be looked at.
  10. DAQ issues
    • CC: CC, Bob, Chanda are working on it (Chandan is leading the effort). We gave a two week timeline before the detectors are pulled out
  11. bmwClient has issues with not pausing when CODA isn't in progress, being stopped and restarted, and to have parameters changed
    • PK will send an email with ideas

Bob's DAQ tasklist

 TO-DO list for Parity DAQ
   updated Jan 4, 2019
Cameron & Bob: 1, 2, 4, 5, 6, 11, 12, 13
Tao & Cameron: 8, 9
Bob: 3, 6, 10
Cip, Cameron, Tao: 13  
1 Doing - Timing: mostly understood, still small mysteries (LNE)
2 Doing - Synchronization-check system. w/help from Ed Jastrzembski
5 To Do - Clean up cabling, labels, fasteners - First need to examine all signals
6 Clean up CRL -- eliminate so-called "dirty trick"
7 Check the 7200 scaler setup at injector, find a spare 7200.
8 Doing - Pedestal studies, aka cross-talk.  
10 Green Monster maintenance.
11 Doing - Test spares.  Repair bad modules.
12 Doing - Document everything.  Circuit diagrams, cable map, "how to".
13 Online software, a bit undefined since software is changing.
    (a) automatic production of run-dependent database
    (b) alarms like end-of-run voice (Bob)
    (c) feedback
    (d) coil-pulsing aka modulation aka dithering (Victoria)
    (e) panguin (GUI showing real-time plots)
    (f) prompt analysis
    (g) e-logbook entries 
    (h) archiving EPICs data


Sanghwa P, Ciprian G, Ye T, Robert R, Amali P, Adam Z, Tao Y, Cameron C, Devi A, Sakib R, Chandan G, Paul K, Kent P


Caryn P